OpenStudy (anonymous):

Solve by completing the square and then use the square root method to solve. x2 - 4x - 32 = 0

OpenStudy (btaylor):

\[x^{2}-4x+4=32+4=36\]\[(x-2)^{2}=36\]\[x-2=\pm \sqrt{36} = \pm 6\]\[x=2 \pm 6\] so x=8 or x=-4

OpenStudy (campbell_st):

start by rewriting the equation to x^2 - 4x = 32 to complete the square you need ( -4/2)^2 = 4 add 4 to both sides x^2 - 4x + 4 = 32 +4 factor the left hand side \[(x - 2)^2 = 36\] take the square root of both sides \[x - 2 = \pm \sqrt{36}\] add 2 to both sides \[x = 2 \pm 6\] x = 8, -4